# See pages - of ArxivIEEE_CopyrightPage_Submission.pdf

### v1

@@ V-F Material comparison
To directly compare the material of the lead glass pieces and the MAVIG BR126 radiation safety glasses, complementary experiments were conducted at Erlangen Centre for Astroparticle Physics (ECAP). The measurements were performed by placing either the radiation safety glasses (MAVIG BR126), lead glass pieces of two different thicknesses (2.5 mm and 2.72 mm) or a large lead glass plate in front of the eye lens dosemeter prototype and irradiating it with radiation qualities of the N-series with a measurement time of 210 s. The measurements were repeated without any shielding to obtain a reference dose. The lead glass plate also had a thickness of 2.5 mm with a surface of about . All materials were placed as close as possible to the eye lens dosemeter prototype. The measured dose quantity in this context is the directional dose equivalent at 3 mm depth as no phantom was included in the setup [8].
Fig. 9 shows the results for irradiations with a collimated and a non-collimated beam. The measured values were normalized to the dose measured without any lead glass shielding. In the case of a non-collimated beam (Fig. 9a), there are significant deviations between the measurements with radiation safety glasses, lead glass pieces and lead glass plate. This is related to the area covered by the material and the resulting influence on the scattering contribution. Since the lead glass pieces only partially shield the eye lens dosemeter prototype, more scattered radiation hits the detector from the side. On the other hand, the lead glass plate covers a large area around the dosemeter prototype resulting in less scattering and lower values. This effect is most pronounced at low mean photon energies as there, the direct beam is almost totally absorbed and, consequently, the detector’s response is dominated by the scattered radiation. At higher energies, the lead glass becomes more transmissive and the differences between the materials decrease.
@@ VI Summary and Conclusion
It has been shown in [9] that the first prototype of the active eye lens dosemeter yields excellent results in measurements of . Those measurements have been performed in compliance with IEC 61526:2024 [10] in front of a cylinder phantom. The results led to interest in whether the dosemeter would also perform well when attached to the side of a phantom, similar to how it is intended to be worn by a final user. The results of this work demonstrate that there is no significant influence on the estimated dose when the measurement is performed at the side of a phantom compared to a position at the center front, for mean photon energies ranging from 12.4 keV to 249 keV as well as for angles between and . However, it has to be pointed out, that these findings were obtained in rather homogeneous radiation fields since the radiation source (X-ray tube) was located at a distance of 2.5 m from the irradiation setup. Thus, corresponding investigations should be repeated in more realistic radiology situations, for instance approximately 50 cm away from a stray body, i.e. a patient, in real life.
@@
Since medical staff usually wear radiation safety glasses when working with ionizing radiation, an attempt was presented to imitate the absorption behavior of such glasses. Small lead glass pieces were attached to the front and the side of the eye lens dosemeter prototype. Additionally, TLDs were installed behind radiation safety glasses for comparison. When irradiated with photons of different angles of incidence, the results of the TLDs showed that the radiation is able to pass through a gap between the nose and the radiation safety glasses at an angle of . For an irradiation angle of , the results of the eye lens dosemeter prototype showed that the radiation bypasses the front lead glass piece and hits the Dosepix detector unshielded. There was a consistent offset between the measured of the TLDs and the eye lens dosemeter prototype, with the prototype measuring lower values. There are several possible explanations for this effect. First, the correction factors used to eliminate the detector differences are energy dependent due to the absorption of the lead glasses affecting the incoming spectra. This effect was considered to be small and was therefore neglected in the analysis. Second, but more significant, is the influence exerted by the lead glass pieces themselves. Comparative measurements were performed to examine the differences between radiation safety glasses and lead glass pieces. The results revealed a strong influence from scattered radiation related to the area shielded by the respective material. The effect was significantly smaller, but still visible with a collimated beam. It is therefore reasonable to assume that these differences between lead glass pieces and radiation safety glasses are responsible for the measured differences of TLDs and eye lens dosemeter prototype. However, while the prototype with lead glass pieces measured lower dose values than the TLDs behind radiation safety glasses, the material comparison showed the opposite behavior. There, the resulting values of the measurements with lead glass pieces were higher than those of the measurements with radiation safety glasses. Ultimately, many different setup properties are possibly influencing the final dose reconstruction. It is still possible that the lead glass pieces are made from a different material since the exact material composition of the lead glass is unknown, or that they have different thicknesses or dimensions. The intercept theorem used for determining the size of the lead glass pieces might not be sufficient. This could be examined by repeating the process with different models of radiation safety glasses. It is also possible that the TLDs measure more backscattered radiation from the phantom. To investigate this effect, a measurement without an Alderson phantom would be advisable. Another possibility is to exchange the EYE-D dosemeter with other passive dosemeters and compare the results with the eye lens dosemeter prototype.
@@
Overall, the positioning of the eye lens dosemeter has proved to be of lower significance, but the reason for the differences between shielding glasses, differences between the eye lens dosemeter prototype and the TLDs as well as the question of which of the two devices comes closest to the actual dose in the eye lens is still unclear. More research is required, but the use of lead glass pieces is a very promising approach.

### v2

@@ V-F Material comparison
To directly compare the material of the lead glass pieces and the MAVIG BR126 radiation safety glasses, complementary experiments were conducted at Erlangen Centre for Astroparticle Physics (ECAP). The eye lens dosemeter prototype was placed at 2 m distance to the radiation source. The measurements were performed by placing either the radiation safety glasses (MAVIG BR126), lead glass pieces of two different thicknesses (2.5 mm and 2.72 mm) or a large, asymmetrical lead glass plate in front of the eye lens dosemeter prototype and irradiating it with radiation qualities of the N-series with a measurement time of 210 s. The measurements were repeated without any shielding to obtain a reference dose. The lead glass plate also had a thickness of 2.5 mm with a width of and height between and . All materials were placed as close as possible to the eye lens dosemeter prototype. The measured dose quantity in this context is the directional dose equivalent at 3 mm depth as no phantom was included in the setup [8].
Fig. 9 shows the results for irradiations with a collimated and a non-collimated beam. The non-collimated beam is used since this represents the realistic conditions in the medical field. For the non-collimated irradiation, the whole setup is illuminated homogeneously. The collimated beam has a diameter of 1 cm. The measured values were normalized to the dose measured without any lead glass shielding. In the case of a non-collimated beam (Fig. 9a), there are significant deviations between the measurements with radiation safety glasses, lead glass pieces and lead glass plate. This is related to the area covered by the material and the resulting influence on the scattering contribution. Since the lead glass pieces only partially shield the eye lens dosemeter prototype, more scattered radiation hits the detector from the side. On the other hand, the lead glass plate covers a large area around the dosemeter prototype resulting in less scattering and lower values. This effect is most pronounced at low mean photon energies as there, the direct beam is almost totally absorbed and, consequently, the detector’s response is dominated by the scattered radiation. At higher energies, the lead glass becomes more transmissive and the differences between the materials decrease.
@@ VI Summary and Conclusion
It has been shown in [9] that the first prototype of the active eye lens dosemeter yields results within the IEC criteria in measurements of . Those measurements have been performed in compliance with IEC 61526:2024 [10] in front of a cylinder phantom. The results led to interest in whether the dosemeter would also perform well when attached to the side of a phantom, similar to how it is intended to be worn by a final user. The results of this work demonstrate that there is no significant influence on the estimated dose when the measurement is performed at the side of a phantom compared to a position at the center front, for mean photon energies ranging from 12.4 keV to 249 keV as well as for angles between and . However, it has to be pointed out, that these findings were obtained in rather homogeneous radiation fields since the radiation source (X-ray tube) was located at a distance of 2.5 m from the irradiation setup. Thus, corresponding investigations should be repeated in more realistic radiology situations, for instance approximately 50 cm away from a stray body, i.e. a patient, in real life.
@@
Since medical staff usually wear radiation safety glasses when working with ionizing radiation, an attempt was presented to imitate the absorption behavior of such glasses. Small lead glass pieces were attached to the front and the side of the eye lens dosemeter prototype. Additionally, TLDs were installed behind radiation safety glasses for comparison. When irradiated with photons of different angles of incidence, the results of the TLDs showed that the radiation is able to pass through a gap between the nose and the radiation safety glasses at an angle of . For an irradiation angle of , the results of the eye lens dosemeter prototype showed that the radiation bypasses the front lead glass piece and hits the Dosepix detector unshielded. There was a consistent offset between the measured of the TLDs and the eye lens dosemeter prototype, with the prototype measuring lower values. There are several possible explanations for this effect. First, the correction factors used to eliminate the detector differences are energy dependent due to the absorption of the lead glasses affecting the incoming spectra. This effect was considered to be small and was therefore neglected in the analysis. Second, but more significant, is the influence exerted by the lead glass pieces themselves. Comparative measurements were performed to examine the differences between radiation safety glasses and lead glass pieces. The results revealed a strong influence from scattered radiation related to the area shielded by the respective material. The effect was significantly smaller, but still visible with a collimated beam. It is therefore reasonable to assume that these differences between lead glass pieces and radiation safety glasses are responsible for the measured differences of TLDs and eye lens dosemeter prototype. However, while the prototype with lead glass pieces measured lower dose values than the TLDs behind radiation safety glasses, the material comparison showed the opposite behavior. There, the resulting values of the measurements with lead glass pieces were higher than those of the measurements with radiation safety glasses. Ultimately, many different setup properties are possibly influencing the final dose reconstruction. It is still possible that the lead glass pieces are made from a different material since the exact material composition of the lead glass is unknown, or that they have different thicknesses or dimensions. The intercept theorem used for determining the size of the lead glass pieces might not be sufficient. This could be examined by repeating the process with different models of radiation safety glasses. It is also possible that the TLDs measure more backscattered radiation from the phantom. Since the rear of the TLDs consists of plastic, similarly to the front, the sensitivity is approximately equal from both sides. On the other hand, it has been shown that the response of eye lens dosemeter prototype increases only by up to 20 % when comparing the measurement position at the side of a water cylinder phantom to the phantom being behind the dosemeter [22]. This indicates that the eye lens dosemeter prototype is less sensitive to scattered radiation than the TLDs. To investigate this effect, a measurement without an Alderson phantom would be advisable. Another possibility is to exchange the EYE-D dosemeter with other passive dosemeters and compare the results with the eye lens dosemeter prototype.
@@
Overall, the positioning of the eye lens dosemeter has proved to be of lower significance, but the reason for the differences between shielding glasses, differences between the eye lens dosemeter prototype and the TLDs as well as the question of which of the two detector types comes closest to the actual dose in the eye lens is still unclear. More research is required, but the use of lead glass pieces is a very promising approach.